OpenStudy (anonymous):

What is the number of ways to arrange 8 objects from a set of 12 different objects?

OpenStudy (anonymous):

You can use here the concept of Combinations..

OpenStudy (jacob902):

12 possible first objects leaves 11 possible second objects leaves 10 possible third objects etc. 12 x 11 x 10 x 9 x 8 x 7 x 6 x 5 = 12! / (12 - 8)!

OpenStudy (anonymous):

well since theres a to choose do 12*11*10*9*8*7*6*5=what @thisotherliz

OpenStudy (anonymous):

You want to just arrange them: So, the number of ways in which \(r\) objects can be arranged from \(n\) different objects is given by: \[\large \implies ^nC_r\]

OpenStudy (anonymous):

what did u get @thisotherliz

OpenStudy (anonymous):

19958400 thank you all

OpenStudy (anonymous):

\[Also \quad \color{green}{^nC_r = \frac{n!}{(n-r)! \cdot r!}}\]

OpenStudy (anonymous):

np dont forget to medal if u need more help just mention me

OpenStudy (anonymous):

I think that is a wrong answer...

OpenStudy (anonymous):

\[^12C_8 = \frac{12!}{5! \times 8!} = \frac{12 \times 11 \times 10 \times 9}{5 \times 4 \times 3 \times 2 \times1} = ??\]

OpenStudy (anonymous):

\[^{12}C_8\]

OpenStudy (anonymous):

I am wrong.. :P

OpenStudy (anonymous):

My mind has got rusty.. :) What the hell yaar..!! Sorry, you all are good.. :)